# Solar Cruiser Disturbance Torque Estimation and Predictive Momentum Management

### v1

@@ 3 Disturbance Estimation Using Kalman Filter
The disturbance torque ${\boldsymbol{\tau}}^{\text{dist}}_{b}$ is an external input acting on the solar sail system that impacts its dynamics. In the MPC policy proposed in Section 4, this disturbance torque is a key parameter in the prediction model used to forecast the system dynamics and determine optimal momentum management actuation. Due to the slow motion and relatively steady attitude operation nature of solar sails, the disturbance torque is modeled as approximately constant or slow varying. However, the disturbance cannot be measured directly, and needs to be estimated onboard. Utilizing a Kalman filter estimation framework provides adequate knowledge of the disturbance that MPC can take advantage of. A Kalman filter framework is developed in this section to supplement this essential parameter for MPC.
@@ 3.1 Measurement Model
$\mbf{z}_{1,k}=\underbrace{\begin{bmatrix}\mbf{1}_{12\times 12}&\mbf{0}_{12\times 3}\end{bmatrix}}_{\mbf{H}_{1}}\begin{bmatrix}\mbf{x}_{k}\\ \mbf{w}_{k}\end{bmatrix}+{\boldsymbol{\nu}}_{1,k},\qquad{\boldsymbol{\nu}}_{1,k}\sim\mathcal{N}(\mbf{0},\mbf{R}^{\text{KF}}_{1}),$
where ${\boldsymbol{\nu}}_{1,k}$ is the linear additive measurement noise, and the measurement error covariance matrix $\mbf{R}^{\text{KF}}_{1}=\text{diag}(\mbf{r}_{\theta},\mbf{r}_{\omega},\mbf{r}_{h},\mbf{r}_{e})$ is determined by the variance of each corresponding state measurement error (${\boldsymbol{\sigma}}_{\theta}^{2},{\boldsymbol{\sigma}}_{\omega}^{2},{\boldsymbol{\sigma}}_{h}^{2},{\boldsymbol{\sigma}}_{e}^{2}$), which is associated with the onboard ADCS state estimation accuracy. Considering that the RW PID control law dictates the reaction wheel angular momentum rate $\dot{\mbf{h}}^{\text{RW}^{\mathsf{T}}}_{b}$, an additional measurement of $\dot{\mbf{h}}^{\text{RW}^{\mathsf{T}}}_{b,k}$ is assumed to be accessible to improve observability, where
$\mbf{z}_{2,k}=\underbrace{\begin{bmatrix}\mbf{K}_{p}&\mbf{K}_{d}&\mbf{0}_{3\times 3}&\mbf{K}_{i}&\mbf{0}_{3\times 3}\end{bmatrix}}_{\mbf{H}_{2}}\begin{bmatrix}\mbf{x}_{k}\\ \mbf{w}_{k}\end{bmatrix}+{\boldsymbol{\nu}}_{2,k},\qquad{\boldsymbol{\nu}}_{2,k}\sim\mathcal{N}(\mbf{0},\mbf{R}^{\text{KF}}_{2}),$
${\boldsymbol{\nu}}_{2,k}$ is the linear additive measurement noise, and the measurement error covariance matrix $\mbf{R}^{\text{KF}}_{2}=\mbf{r}_{\dot{h}}$ is associated to the variance of each corresponding state measurement error in $\mbf{R}^{\text{KF}}_{1}$.
The complete Kalman measurement model is formulated as
$\underbrace{\begin{bmatrix}\mbf{z}_{1,k}\\ \mbf{z}_{2,k}\end{bmatrix}}_{\mbf{Z}_{k}}=\underbrace{\begin{bmatrix}\mbf{H}_{1}\\ \mbf{H}_{2}\end{bmatrix}}_{\mbf{H}}\begin{bmatrix}\mbf{x}_{k}\\ \mbf{w}_{k}\end{bmatrix}+\underbrace{\begin{bmatrix}{\boldsymbol{\nu}}_{1,k}\\ {\boldsymbol{\nu}}_{2,k}\end{bmatrix}}_{{\boldsymbol{\nu}}_{k}},\quad{\boldsymbol{\nu}}_{k}\sim\mathcal{N}(\mbf{0},\mbf{R}^{\text{KF}}),$
where $\mbf{R}^{\text{KF}}=\text{diag}(\mbf{r}_{\theta},\mbf{r}_{\omega},\mbf{r}_{h},\mbf{r}_{e},\mbf{r}_{\dot{h}})$.
@@ 3.2 Process Model
Given the slow evolution of the spacecraft attitude, it is assumed that SRP force $\mbf{f}_{b}^{\text{SRP}}$ is a known constant, the error of the dynamic model is Gaussian and linearly additive, and the disturbance torque to be estimated, ${\boldsymbol{\tau}}_{b}^{\text{dist}}$, is constant. The discrete-time LTV model in Eq. (10) with the addition of linear model error is given by
$\mbf{x}^{\text{KF}}_{k+1}$ $=\mbf{A}_{k}\mbf{x}^{\text{KF}}_{k}+\mbf{B}_{w,k}\mbf{w}^{\text{KF}}_{k}+\mbf{B}_{u1,k}\mbf{u}^{\text{AMT}}_{k}+\mbf{B}_{u2,k}{u}^{\text{RCD}}_{k}+{\boldsymbol{\eta}}^{\text{model}}_{k},\qquad{\boldsymbol{\eta}}^{\text{model}}_{k}\sim\mathcal{N}(\mbf{0},\mbf{Q}^{\text{KF}}_{\text{model}}),$
@@
where the linear model error ${\boldsymbol{\eta}}^{\text{model}}_{k}$ and disturbance error ${\boldsymbol{\eta}}^{\text{dist}}_{k}$ are assumed to be linearly additive and normally distributed. The error covariance matrices of the model uncertainty are $\mbf{Q}^{\text{KF}}_{\text{model}}=\text{diag}(\mbf{q}_{\theta},\mbf{q}_{\omega},\mbf{q}_{h},\mbf{q}_{e})$ and $\mbf{Q}^{\text{KF}}_{\text{dist}}=\text{diag}({q}_{\tau 1},{q}_{\tau 2},{q}_{\tau 3})$, which are tuning parameters chosen to influence the Kalman filter’s aggressiveness in updating the estimate of the disturbance torque $\mbf{w}^{\text{KF}}$. In particular, increasing the variances within $\mbf{Q}^{\text{KF}}_{\text{model}}$ results in a slower convergence of the disturbance torque, while increase the variances within $\mbf{Q}^{\text{KF}}_{\text{dist}}$ speeds up the convergence, while potentially making the estimates more sensitive to measurement noise.
@@
$\underbrace{\begin{bmatrix}\hat{\mbf{x}}^{-}_{k+1}\\ \hat{\mbf{w}}^{-}_{k+1}\end{bmatrix}}_{\hat{\mbf{X}}^{-}_{k+1}}=\underbrace{\begin{bmatrix}\mbf{A}_{k}&\mbf{B}_{w,k}\\ \mbf{0}_{3\times 12}&\mbf{1}_{3\times 3}\end{bmatrix}}_{\mbf{F}_{k}}\underbrace{\begin{bmatrix}\hat{\mbf{x}}^{+}_{k}\\ \hat{\mbf{w}}^{+}_{k}\end{bmatrix}}_{\hat{\mbf{X}}^{+}_{k}}+\underbrace{\begin{bmatrix}\mbf{B}_{u,k}\\ \mbf{0}_{3\times 3}\end{bmatrix}}_{\mbf{G}_{k}}\underbrace{\begin{bmatrix}\mbf{u}^{\text{AMT}}_{k}\\ {u}^{\text{RCD}}_{k}\end{bmatrix}}_{\mbf{U}_{k}}+\underbrace{\begin{bmatrix}{\boldsymbol{\eta}}^{\text{model}}_{k}\\ {\boldsymbol{\eta}}^{\text{dist}}_{k}\end{bmatrix}}_{{\boldsymbol{\eta}}_{k}},$
@@ 4.2 Prediction Model
$\dot{\mbf{x}}^{\text{MPC}}=\mbf{A}^{\text{MPC}}\mbf{x}^{\text{MPC}}+\mbf{B}_{w}^{\text{MPC}}\mbf{w}+\mbf{B}_{u1}^{\text{MPC}}\mbf{u}^{\text{AMT}}+\mbf{B}_{u2}^{\text{MPC}}u^{\text{RCD}},$ (11)
where $\mbf{A}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{A}\mbf{T}$, $\mbf{B}_{w}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{B}_{w}$, $\mbf{B}_{u1}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{B}_{u1}$, and $\mbf{B}_{u2}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{B}_{u2}$. A ZOH is then applied to the inputs of Eq. (11) to yield the discrete-time linear prediction used by MPC over its prediction model, given by
$\mbf{x}^{\text{MPC}}_{j+1|t_{k}}=\mbf{A}^{\text{MPC}}_{k}\mbf{x}^{\text{MPC}}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{w,k}\mbf{w}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{u1,k}\mbf{u}^{\text{AMT}}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{u2,k}{u}^{\text{RCD}}_{j|t_{k}},\qquad j=0,1,\ldots,N-1,$ (12)
where the subscript $j|t_{k}$ refers to the $j$-th discrete time step within the MPC prediction horizon at time step $t_{k}$. The LTV matrices of $\mbf{A}^{\text{MPC}}_{k}$, $\mbf{B}^{\text{MPC}}_{w,k}$, $\mbf{B}^{\text{MPC}}_{u1,k}$, $\mbf{B}^{\text{MPC}}_{u2,k}$ are updated at every momentum management timestep at time $t_{k}$, and kept constant throughout the MPC prediction horizon, resulting in a linear time-invariant (LTI) model over the span of the prediction horizon. The MPC framework uses the state estimate from Kalman filter framework presented in Section 3 as its knowledge within the prediction model, where $\mbf{x}^{\text{MPC}}_{0|t_{k}}=\begin{bmatrix}\hat{{\boldsymbol{\theta}}}_{k}^{\mathsf{T}}&\hat{{\boldsymbol{\omega}}}^{ba^{\mathsf{T}}}_{b,k}&\mathbf{M}_{34}^{\dagger}\hat{\mbf{h}}_{b,k}^{\text{RW}^{\mathsf{T}}}&\hat{\mbf{e}}_{k}^{\text{int}^{\mathsf{T}}}\end{bmatrix}^{\mathsf{T}}$ and $\mbf{w}_{j|t_{k}}=\hat{\mbf{w}}^{+}_{k}$, $j=0,1,\ldots,N-1$.
@@ 4.3 State Constraints
To ensure the practical feasibility of the controller, inequality constraints are imposed on the system states. These constraints enforce bounded deviations in sailcraft attitude, angular velocity, RW angular momentum, and integrated attitude error across the prediction horizon as ${\boldsymbol{\theta}}_{\text{min}}\leq{\boldsymbol{\theta}}_{j|t_{k}}\leq{\boldsymbol{\theta}}_{\text{max}}$, ${\boldsymbol{\omega}}^{ba}_{b,\text{min}}\leq{\boldsymbol{\omega}}_{b,j|t_{k}}^{ba}\leq{\boldsymbol{\omega}}^{ba}_{b,\text{max}}$, $\mbf{h}^{\text{RW}}_{4,\text{min}}\leq\mbf{h}_{4,j|t_{k}}^{\text{RW}}\leq\mbf{h}^{\text{RW}}_{4,\text{max}}$, and $\mbf{e}^{\text{int}}_{\text{min}}\leq\mbf{e}^{\text{int}}_{j|t_{k}}\leq\mbf{e}^{\text{int}}_{\text{max}}$, respectively. Collectively, this is written as the constraint $\mbf{x}^{\text{MPC}}_{\text{min}}\leq\mbf{x}^{\text{MPC}}_{j|t_{k}}\leq\mbf{x}^{\text{MPC}}_{\text{max}}$.

### v2

@@ 3 Disturbance Estimation Using Kalman Filter
Given NASA’s efforts to develop on-orbit SRP calibration algorithms (Carzana et al., 2023), it is reasonable to assume that an accurate SRP force estimate is available within the proposed momentum management algorithm. However, the SRP disturbance torque cannot be measured directly, and needs to be estimated onboard in real time. The disturbance torque ${\boldsymbol{\tau}}^{\text{dist}}_{b}$ is an external input acting on the solar sail system that impacts its dynamics. In the MPC policy proposed in Section 4, this disturbance torque is a key parameter in the prediction model used to forecast the system dynamics and determine optimal momentum management actuation. Due to the slow motion and relatively steady attitude operation nature of solar sails, the disturbance torque is modeled as approximately constant or slow varying. A Kalman filter framework is developed in this section to supplement this essential parameter for MPC momentum management.
@@ 3.1 Measurement Model
$\mbf{y}_{k}=\underbrace{\begin{bmatrix}\mbf{1}_{12\times 12}&\mbf{0}_{12\times 3}\end{bmatrix}}_{\mbf{H}}\begin{bmatrix}\mbf{x}_{k}\\ \mbf{w}_{k}\end{bmatrix}+{\boldsymbol{\nu}}_{k},\qquad{\boldsymbol{\nu}}_{k}\sim\mathcal{N}(\mbf{0},\mbf{R}^{\text{KF}}),$
where ${\boldsymbol{\nu}}_{k}$ is the linear additive measurement noise, and the measurement error covariance matrix $\mbf{R}^{\text{KF}}=\text{diag}(\mbf{r}_{\theta},\mbf{r}_{\omega},\mbf{r}_{h},\mbf{r}_{e})$ is determined by the variance of each corresponding state measurement error (${\boldsymbol{\sigma}}_{\theta}^{2},{\boldsymbol{\sigma}}_{\omega}^{2},{\boldsymbol{\sigma}}_{h}^{2},{\boldsymbol{\sigma}}_{e}^{2}$), which is associated with the onboard ADCS state estimation accuracy.
@@ 3.2 Process Model
Given the slow evolution of the spacecraft attitude, it is assumed that SRP force ${\bar{\mbf{f}}}_{b}^{\text{SRP}}$ is a known constant updated at every time step, the error of the dynamics model is Gaussian and linearly additive, and the disturbance torque to be estimated, ${\boldsymbol{\tau}}_{b}^{\text{dist}}$, is constant within the time update (prediction) step. The discrete-time LTV model in Eq. (10) with the addition of linear model error is given by
$\mbf{x}^{\text{KF}}_{k+1}$ $=\mbf{A}_{k}\mbf{x}^{\text{KF}}_{k}+\mbf{B}_{w,k}\mbf{w}^{\text{KF}}_{k}+\mbf{B}_{u1,k}\mbf{u}^{\text{AMT}}_{k}+\mbf{B}_{u2,k}{u}^{\text{RCD}}_{k}+\mbf{z}_{k}+{\boldsymbol{\eta}}^{\text{model}}_{k},\qquad{\boldsymbol{\eta}}^{\text{model}}_{k}\sim\mathcal{N}(\mbf{0},\mbf{Q}^{\text{KF}}_{\text{model}}),$
@@
where the linear model error ${\boldsymbol{\eta}}^{\text{model}}_{k}$ and disturbance error ${\boldsymbol{\eta}}^{\text{dist}}_{k}$ are assumed to be linearly additive and normally distributed. The error covariance matrices of the model uncertainty are $\mbf{Q}^{\text{KF}}_{\text{model}}=\text{diag}(\mbf{q}_{\theta},\mbf{q}_{\omega},\mbf{q}_{h},\mbf{q}_{e})$ and $\mbf{Q}^{\text{KF}}_{\text{dist}}$, which are tuning parameters chosen to influence the Kalman filter’s aggressiveness in updating the estimate of the disturbance torque $\mbf{w}^{\text{KF}}$. In particular, increasing the variances within $\mbf{Q}^{\text{KF}}_{\text{model}}$ results in a slower convergence of the disturbance torque, while increase the variances within $\mbf{Q}^{\text{KF}}_{\text{dist}}$ speeds up the convergence, while potentially making the estimates more sensitive to measurement noise. In order to adapt to the attitude-dependent disturbance torque varying with the slew maneuver, a dynamic error covariance scaling quadratically with the desired angular velocity is added to the disturbance covariance matrix, such that $\mbf{Q}^{\text{KF}}_{\text{dist}}=\text{diag}({q}_{\tau,1}+\xi_{1}\omega_{d,1}^{2},{q}_{\tau,2}+\xi_{2}\omega_{d,2}^{2},{q}_{\tau,3}+\xi_{3}\omega_{d,3}^{2})$, and $\xi_{i}$ is the scaling parameter associate to the dynamic covariance for $i=1,2,3$.
@@
$\underbrace{\begin{bmatrix}\hat{\mbf{x}}^{-}_{k+1}\\ \hat{\mbf{w}}^{-}_{k+1}\end{bmatrix}}_{\hat{\mbf{X}}^{-}_{k+1}}=\underbrace{\begin{bmatrix}\mbf{A}_{k}&\mbf{B}_{w,k}\\ \mbf{0}_{3\times 12}&\mbf{1}_{3\times 3}\end{bmatrix}}_{\mbf{F}_{k}}\underbrace{\begin{bmatrix}\hat{\mbf{x}}^{+}_{k}\\ \hat{\mbf{w}}^{+}_{k}\end{bmatrix}}_{\hat{\mbf{X}}^{+}_{k}}+\underbrace{\begin{bmatrix}\mbf{B}_{u,k}\\ \mbf{0}_{3\times 3}\end{bmatrix}}_{\mbf{G}_{k}}\underbrace{\begin{bmatrix}\mbf{u}^{\text{AMT}}_{k}\\ {u}^{\text{RCD}}_{k}\end{bmatrix}}_{\mbf{U}_{k}}+\underbrace{\begin{bmatrix}\mbf{z}_{k}\\ \mbf{0}\end{bmatrix}}_{\mbf{Z}_{k}}+\underbrace{\begin{bmatrix}{\boldsymbol{\eta}}^{\text{model}}_{k}\\ {\boldsymbol{\eta}}^{\text{dist}}_{k}\end{bmatrix}}_{{\boldsymbol{\eta}}_{k}},$
@@ 4.1 Introduction to MPC
4.2 RCD Constraint Relaxation
The control input $u^{\text{RCD}}=\tau^{\text{RCD}}_{b3}$ is chosen based on the momentum management strategy. Considering the RCD on-off actuation as an explicit constraint in the MPC optimization problem leads to a mixed-integer problem, which is computationally expensive, and limits the practicality of onboard real-time MPC. To enable the use of convex optimization solvers with the proposed MPC policy, the integer constraint is relaxed, and a PWM quantization is applied to the RCD actuation (Shen and Caverly, 2026). A continuous value of $u^{\text{RCD}}_{\text{mpc}}$ is allowed in the optimization problem, where $\tau^{\text{RCD}}_{b3,\text{on}}\leq u^{\text{RCD}}_{\text{mpc}}\leq\tau^{\text{RCD}}_{b3,\text{on}}$ and $\tau^{\text{RCD}}_{b3,\text{on}}$ is the roll torque magnitude generated when the RCDs are turned on. After solving the MPC optimization problem, the continuous $u^{\text{RCD}}_{\text{mpc}}$ is then quantized into a discrete value
where $\beta_{\text{on}}\in\{-1,1\}$ denotes the clockwise and counterclockwise directions about the roll ($\underrightarrow{b}^{3}$) axis, $\tau^{\text{RCD}}_{b3,\text{on}}$ denotes the torque magnitude when RCDs are turned on, and $t_{c}=\Delta t\cdot\frac{u^{\text{RCD}}_{\text{mpc}}}{\tau^{\text{RCD}}_{b3,\text{on}}}$ is the length (cut-off time) of a single pulse PWM conversion from a continuous MPC optimal RCD input. Details of the RCD quantization can be found in the work of Shen and Caverly (2026).
@@ 4.3 Prediction Model
$\dot{\mbf{x}}^{\text{MPC}}=\mbf{A}^{\text{MPC}}\mbf{x}^{\text{MPC}}+\mbf{B}_{w}^{\text{MPC}}\mbf{w}+\mbf{B}_{u1}^{\text{MPC}}\mbf{u}^{\text{AMT}}+\mbf{B}_{u2}^{\text{MPC}}u^{\text{RCD}}+\mbf{z}^{\text{MPC}},$ (12)
where $\mbf{A}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{A}\mbf{T}$, $\mbf{B}_{w}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{B}_{w}$, $\mbf{B}_{u1}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{B}_{u1}$, $\mbf{B}_{u2}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{B}_{u2}$, and $\mbf{z}^{\text{MPC}}=\mbf{T}^{\dagger}\mbf{z}$. A ZOH is then applied to the inputs of Eq. (12) to yield the discrete-time linear prediction used by MPC over its prediction model, given by
$\mbf{x}^{\text{MPC}}_{j+1|t_{k}}=\mbf{A}^{\text{MPC}}_{j|t_{k}}\mbf{x}^{\text{MPC}}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{w,j|t_{k}}\mbf{w}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{u1,j|t_{k}}\mbf{u}^{\text{AMT}}_{j|t_{k}}+\mbf{B}^{\text{MPC}}_{u2,j|t_{k}}{u}^{\text{RCD}}_{j|t_{k}}+\mbf{z}^{\text{MPC}}_{j|t_{k}},\qquad j=0,1,\ldots,N-1,$ (13)
where the subscript $j|t_{k}$ refers to the $j$-th discrete time step within the MPC prediction horizon at time step $t_{k}$. The prediction model includes the state $\mbf{x}^{\text{MPC}}_{j|t_{k}}=\begin{bmatrix}{\boldsymbol{\theta}}_{j|t_{k}}^{\mathsf{T}}&{\boldsymbol{\omega}}^{ba^{\mathsf{T}}}_{b,j|t_{k}}&\mbf{h}_{4,j|t_{k}}^{\text{RW}^{\mathsf{T}}}&\mbf{e}_{j|t_{k}}^{\text{int}^{\mathsf{T}}}\end{bmatrix}^{\mathsf{T}}$, the AMT input $\mbf{u}^{\text{AMT}}_{j|t_{k}}$, and the RCD input ${u}^{\text{RCD}}_{j|t_{k}}$ to be designed across the prediction horizon $j=0,1,\ldots,N-1$ at time $t_{k}$. The disturbance torque is represented by $\mbf{w}_{j|t_{k}}$. The MPC framework uses the state estimate from Kalman filter framework presented in Section 3 as its knowledge within the prediction model, where $\mbf{x}^{\text{MPC}}_{0|t_{k}}=\begin{bmatrix}\hat{{\boldsymbol{\theta}}}_{k}^{\mathsf{T}}&\hat{{\boldsymbol{\omega}}}^{ba^{\mathsf{T}}}_{b,k}&\mathbf{M}_{34}^{\dagger}\hat{\mbf{h}}_{b,k}^{\text{RW}^{\mathsf{T}}}&\hat{\mbf{e}}_{k}^{\text{int}^{\mathsf{T}}}\end{bmatrix}^{\mathsf{T}}$ and $\mbf{w}_{j|t_{k}}=\hat{\mbf{w}}^{+}_{k}$ for $j=0,1,\ldots,N-1$. The $\hat{(\cdot)}^{+}$ notation for a posteriori (measurement updated) state estimate is simplified as $\hat{(\cdot)}$ to avoid notation clustering. The LTV matrices of $\mbf{A}^{\text{MPC}}_{j|t_{k}}$, $\mbf{B}^{\text{MPC}}_{w,j|t_{k}}$, $\mbf{B}^{\text{MPC}}_{u1,j|t_{k}}$, $\mbf{B}^{\text{MPC}}_{u2,j|t_{k}}$, $\mbf{z}^{\text{MPC}}_{j|t_{k}}$ are linearized about the operation points (varying with the slew maneuver) across the prediction horizon, where $\bar{\mbf{x}}^{\text{MPC}}_{j|t_{k}}=\Big[{\boldsymbol{\theta}}_{d,j|t_{k}}^{\mathsf{T}}\,\,\,{\boldsymbol{\omega}}_{d,j|t_{k}}^{{\mathsf{T}}}\,\,\,\mathbf{M}_{34}^{\dagger}\hat{\mbf{h}}_{b,k}^{\text{RW}^{\mathsf{T}}}\,\,\,\hat{\mbf{e}}^{\text{int}^{\mathsf{T}}}_{k}\Big]^{\mathsf{T}}$ for $j=0,1,\ldots,N-1$, ${\bar{\mbf{r}}}^{ps}_{b}={\bar{\mbf{r}}}^{ps}_{b}(t_{k})$, $\bar{\mbf{f}}_{b}^{\text{SRP}}=\mbf{f}_{b}^{\text{SRP}}(t_{k})$ are updated at every momentum management timestep at time $t_{k}$, and $\dot{\bar{\mbf{r}}}_{b}^{ps}=\ddot{\bar{\mbf{r}}}_{b}^{ps}=\mbf{0}$, $\bar{{\boldsymbol{\tau}}}^{\text{RCD}}_{b}=\mbf{0}$. The time-varying parameters are the desired attitude ${\boldsymbol{\theta}}_{d,j|t_{k}}$ and desired angular velocity ${\boldsymbol{\omega}}_{d,j|t_{k}}$, while the SRP force, disturbance torque, state (angular momentum and integral state), and AMT position to be linearized about are kept constant throughout the MPC prediction horizon.
@@ 4.4 State Constraints
To ensure the practical feasibility of the controller, inequality constraints are imposed on the system states. These constraints enforce bounded deviations in sailcraft attitude, angular velocity, RW angular momentum, and integrated attitude error across the prediction horizon as ${\boldsymbol{\theta}}_{d,j|t_{k}}-{\boldsymbol{\theta}}_{\text{err}}\leq{\boldsymbol{\theta}}_{j|t_{k}}\leq{\boldsymbol{\theta}}_{d,j|t_{k}}+{\boldsymbol{\theta}}_{\text{err}}$, ${\boldsymbol{\omega}}_{d,j|t_{k}}-{\boldsymbol{\omega}}_{\text{err}}\leq{\boldsymbol{\omega}}_{b,j|t_{k}}^{ba}\leq{\boldsymbol{\omega}}_{d,j|t_{k}}+{\boldsymbol{\omega}}_{\text{err}}$, $\mbf{h}^{\text{RW}}_{4,\text{min}}\leq\mbf{h}_{4,j|t_{k}}^{\text{RW}}\leq\mbf{h}^{\text{RW}}_{4,\text{max}}$, and $\mbf{e}^{\text{int}}_{\text{min}}\leq\mbf{e}^{\text{int}}_{j|t_{k}}\leq\mbf{e}^{\text{int}}_{\text{max}}$, respectively. The variables ${\boldsymbol{\theta}}_{\text{err}}$ and ${\boldsymbol{\omega}}_{\text{err}}$ are the allowable deviation from the nominal slew trajectory ${\boldsymbol{\theta}}_{d,j|t_{k}}$ and ${\boldsymbol{\omega}}_{d,j|t_{k}}$ across the prediction horizon. Collectively, this is written as the constraint $\mbf{x}^{\text{MPC}}_{\text{min}}\leq\mbf{x}^{\text{MPC}}_{j|t_{k}}\leq\mbf{x}^{\text{MPC}}_{\text{max}}$.